1x+6y=372 and 4x+12=780 solve system using elimination
Answer by mananth(16946)

1 x + 6 y = 372 .............1

4 x + 12 y = 780 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-2 x -12 y = -744
4 x + 12 y = 780
Add the two equations
2 x = 36
/ 2
x = 18
plug value of x in (1)
1 x + 6 y = 372
18 + 6 y = 372
6 y = 372 -18
6 y = 354
y = 59
x= 18
y= 59
m.ananth@hotmail.ca